find the domain and range of f(x)=x/2-3x
if 2i+5j+k and 4i+mj+nkare collinear vectors then find value of x

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
f(x) = x/(2-3x)
Domain = R - {2/3}
 
For range
y = x/(2-3x)
2y - 3xy = x
x = 2y/(1+3y)
 
Hence
Range = R - {-1/3}
 
 
Regards
Arun (askIITians forum expert)